Find the inverse, f-1(x), for each function 7. f(x) x3 2х+3 8. f (x) 5x4
Answer #1

7 for = kt3 - 0 Let y= f(r) Putin g= ret3 swap re withy re=y+3 solve for y y=43 je g = f cre) (bez randy swap) se tcre) = xf Cre) = 2x+3 - 0 5k + 4 het y = f(x) putiting in o y= 26t3 56 + 4 swap x with y x = 24 +3 54+4 Sky +44 24+3 (5r-2) Y = 3-4K